OpenStudy (anonymous):

an escalator at a shopping center is 200 ft 9 in. long, and rises at an angle of 15 degrees.What is the vertical rise of the escalator?round to the nearest inch

OpenStudy (campbell_st):

draw a picture but first change 200ft 9 in to feet so its 200.75 ft  now you have a right triangle that has a hypotenuse and you need to find the opposite side... which trig ratio is needed..?

OpenStudy (anonymous):

im not sure is it the sin?

OpenStudy (campbell_st):

thats it... so you have \[\sin(15) = \frac{h}{200.75}\] noe solve for h, the vertical height

OpenStudy (anonymous):

idk how to solve for h

OpenStudy (campbell_st):

multiply both sides of the equation by 200.75 so you have \[200.75 \times \sin(15) = h\] noe just do the calculation.

OpenStudy (anonymous):

is it 51.960511?

OpenStudy (campbell_st):

yes... thats it.. so I'm not sure as to the accuracy of the measurement... it could be 52 ft of 51.96 ft

OpenStudy (anonymous):

thank you veryyyy much :)